# Dataset of GenAI-Assisted Information Problem Solving in Education

### v1

@@ Abstract
Information Problem Solving (IPS) is a critical competency for academic and professional success in education, work, and life. The advent of Generative Artificial Intelligence (GenAI), particularly tools like ChatGPT, has introduced new possibilities for supporting students in complex IPS tasks. However, empirical insights into how students engage with GenAI during IPS and how these tools can be effectively leveraged for learning remain limited. Moreover, differences in background—shaped by cultural and socioeconomic factors—pose additional challenges to the equitable integration of GenAI in educational contexts. To address this gap, we present an open-source dataset collected from 279 students at a public Australian university. The dataset was generated through students’ use of FLoRA, a GenAI-powered educational platform that widely adopted in the field of learning analytics. Within FLoRA, students interacted with an embedded GenAI chatbot to gather information and synthesize it into data science project proposals. The dataset captures fine-grained, multi-dimensional records of GenAI-assisted IPS processes, including: (i) student–GenAI dialogue transcripts; (ii) writing process log traces; (iii) final project proposals with human-assigned assessment scores; (iv) surveys of biographic and prior knowledge in data science and AI; and (v) surveys capturing students’ GenAI experience and perceptions of GenAI’s effectiveness in supporting IPS. This dataset provides a valuable resource for advancing our understanding of GenAI’s role in educational IPS and informing the design of adaptive, inclusive AI-powered learning tools.
@@
In recent years, GenAI systems have been rapidly adopted by learners to support a wide range of educational activities, including IPS tasks [10]. Powered by large language models, GenAI tools enable learners to interact with information in conversational and iterative ways, thereby extending traditional forms of tool use in IPS. Rather than being confined to a single stage, GenAI can be integrated across multiple phases of the IPS process [11, 12]. During the early stages of IPS, learners may use GenAI to clarify task requirements, refine research questions, or decompose complex information problems into manageable sub-tasks. Beyond problem formulation, GenAI increasingly serves as an interactive channel for information gathering, allowing learners to obtain explanations, examples, and summaries through iterative dialogue rather than through conventional keyword-based search alone. In later phases, learners may rely on GenAI to summarise retrieved materials, compare perspectives across sources, generate integrative explanations, and obtain feedback on intermediate or final task products (e.g., essay drafts) to support evaluation and revision. Collectively, these practices indicate that GenAI functions as a flexible cognitive tool embedded throughout learners’ IPS workflows, with implications for both the nature and sequencing of IPS activities.
@@
The rapid integration of GenAI into students’ learning practices has raised important questions about its potential impact on learning processes and outcomes. While GenAI offers new opportunities to support information access, sense-making, and feedback [13], concerns have also been expressed regarding over-reliance, reduced cognitive and metacognitive engagement, and exposure to inaccuracies and biases that may compromise information quality [14, 15, 16, 17]. These competing possibilities underscore the need for systematic empirical investigation into how learners actually make use of GenAI in their everyday IPS tasks. Although a growing body of research has begun to examine students’ use of GenAI in educational contexts, existing studies have often focused on isolated IPS phases, relied on controlled laboratory settings, or drawn primarily on learners’ retrospective self-reports, thereby providing only partial and indirect insights into real-world IPS practices. To address these limitations, we argue that a critical next step is to collect and openly share large-scale datasets that capture learners’ fine-grained behavioural traces as they engage in IPS processes with GenAI in authentic educational scenarios. Such datasets are essential for advancing data-driven understanding of GenAI-assisted IPS and for informing evidence-based instructional design and educational policy.
@@
To address these challenges, this paper presents a large-scale, openly available dataset capturing learners’ IPS activities in an authentic GenAI-assisted educational context. The dataset was collected through a field study involving 519 postgraduate students at a public Australian university, who completed a two-week course assignment to develop a data science project proposal. Students were supported by FLoRA, a GenAI-powered system that integrates a GPT-4o chatbot with note-taking and writing tools and is designed to provide scaffolded support rather than direct solutions. The dataset comprises multiple complementary data modalities, including student–GenAI dialogue logs, fine-grained interaction traces of the writing process (e.g., mouse movements, keystrokes, and clicks), the project proposals produced as task outcomes together with instructor-assigned evaluation scores, and survey data capturing learners’ background characteristics and perceptions of GenAI use. By capturing learners’ interactions with GenAI as they gather information and construct project proposals in an authentic course setting, this dataset provides a unique empirical foundation for investigating IPS processes across phases and for developing data-driven instructional and AI-based scaffolding approaches to support effective student learning.
@@ Methods
This research project rigorously adheres to the ethical standards established by Monash University. The study was approved by Monash Human Research Ethics Committee under Project ID 42510. The dataset was collected from students in developing a data science project proposal as an assignment in an introductory level data science course for postgraduate students. Before commencing data collection, we distributed explanatory statements and forum announcements to all students to introduce the purpose of the study and to clarify that the data collected would be used exclusively for research purposes. Given that the study requires the use of FLoRA [18], we informed students that participation was optional and that they could complete the assignment without utilizing FLoRA if they preferred. We emphasised that all students had the right to withdraw from the research study at any time. Importantly, we assured them that their decision to participate or not would have no influence on their final assignment scores. By providing this information, we ensured that students could make informed decisions about their involvement, thereby upholding ethical research practices and respecting their autonomy.
@@
The participants in this study comprised 519 postgraduate students and 65 students choose not to use FLoRA. These students had enrolled in the Foundations of Data Science course, either as an elective or a core component of their respective degrees. As this research was conducted as a field study, all participants worked within the same learning environment, ensuring consistency in the educational context.
@@
Data collection was conducted using the FLoRA platform [18], which offers a suite of tools tailored for educational research, including an annotation tool, an AI-powered chatbot tool, and a writing tool. The study was systematically divided into four key phases: Pre-task survey, Training task, Main task, and Post-task survey. Each phase was designed to collect specific types of data. By capturing fine-grained, time-stamped logs of students’ annotations, trace of using the platform and interactions with the chatbot in real time, FLoRA yields high-resolution, process-oriented data on learners’ information problem-solving activities.
@@ 3. Phase 3:
    GenAI-Assisted Project Proposal Task. In Phase 3, students were required to develop a data science project proposal using the FLoRA platform, supported by an embedded GenAI-powered chatbot. They were first asked to select one topic from a set of 20 options spanning domains such as agriculture, education, finance, the game industry, healthcare, social media, and sports. Based on the chosen topic, students then had to formulate an idea for how data science could be applied in that domain and elaborate it into a written project proposal. Throughout this process, the chatbot functioned as an interactive research assistant, enabling students to ask questions, seek clarification, and obtain relevant information to support their information problem-solving activities. For example, a student exploring a topic on urban mobility might ask, “What are common data sources used in transportation studies?” or “Can you explain the concept of traffic flow prediction using machine learning?”. Also, students might use the chatbot to refine their ideas, test the relevance of their project scope, or request feedback on parts of their written proposal, such as asking, “Does this hypothesis seem focused enough?” or “Can you suggest improvements to this paragraph?” While the chatbot offered rich, responsive support, it was deliberately constrained from generating complete proposal content to ensure that students engaged meaningfully with the task and developed their own academic output. In this phase, trace data including the annotations students made, the interactions with the platform and the writing process logs was also recorded.
@@ 4. Phase 4:
  • Missing writing trace data: Students who used external tools (e.g., Microsoft Word or Google Docs) to compose their proposals instead of FLoRA’s built-in writing tool were removed, since writing logs, which are crucial for tracing IPS-related behaviours, could not be captured.
@@
  • Low similarity between FLoRA and Moodle submissions: Students were required to copy their FLoRA-composed proposal into a Microsoft Word document and submit it via Moodle for grading. To ensure consistency between the analysed proposal in FLoRA and the final assessed work submitted on Moodle, we computed similarity scores between both versions. Proposals with a similarity score below 0.85 (this threshold of 0.85 was determined through manual review of the lowest 10% most dissimilar cases) were excluded due to substantial post-transfer modifications.
@@
We began with the taxonomy proposed in the paper of Ren et al. [20] but found that an excessive number of student utterances were being assigned to the general-purpose category Reveal (i.e., expressing a new intent for information or refining a previous one). To enhance analytical granularity and improve interpretability, we renamed Reveal as Idea Exploration and subdivided it into six distinct student intent types: Knowledge, Judgement, Preference, Provide Context, Revise, and Inspire. The first three subcategories—Knowledge, Judgement, and Preference—were informed by Paul and Elder’s critical thinking framework [21], which distinguishes between factual, evaluative, and subjective forms of reasoning. Specifically, Knowledge utterances involve requests for definitive, factual information that is consistent across contexts (e.g., definitions or technical terms). Judgement utterances seek responses that involve critical evaluation or consideration of multiple valid perspectives. Preference utterances solicit subjective opinions or recommendations, such as which idea might be easier or more appropriate. The remaining three subcategories—Provide Context, Revise, and Inspire—emerged inductively through close examination of our collected data. Provide Context captures utterances in which students offer background information or setup without making a direct information request, often to improve the relevance of future responses. Revise applies to instances where students reformulate or elaborate on a previous question, typically in response to a chatbot answer that was unclear or insufficient. Inspire is used when students pose follow-up questions that are clearly triggered by the chatbot’s preceding response, indicating a sustained, dynamic information-seeking process.
@@
All dialogue data were annotated by three independent coders with relevant expertise in data science and educational technology. To ensure clarity and consistency in the annotation process, a detailed coding manual was developed, including operational definitions, decision rules, and illustrative examples for each code. This preparatory work was essential for aligning coders’ interpretations and reducing ambiguity during the annotation of both student intents and chatbot responses.
@@
In total, the annotated dataset comprises 3,560 student prompts and chatbot responses, each labelled according to the developed codebooks described in Table 1 and Table 2, with the distribution of intent and response types visualised in Figure 1 and the distribution student activity visualised in Figure 2.
@@
The dataset is publicly available on GitHub [^1] https://github.com/Xinyu-Li/FLoRA_Open_Data_Monash_FIT5145 under the BSD-2-Clause license. It comprises multiple CSV-format data files, as detailed below.
@@
Table 3: The learning action categories
No. Action No. Sub_Action Sub Action Definition
0 START 0.1 START_TASK Indicate when learners start the task.
2 READING 2.1 READING_MATERIALS Learners re-read relevant reading materials.
2.2 REREADING_MATERIALS Learners re-read relevant reading materials.
3.4 SAVE_ESSAY Learners click the save button to save the essay.
3.5 CLOSE_ESSAY Learners close the essay window.
4.8 CLOSE_ANNOTATION Learners close the annotation window.
5 CHATBOT 5.1 OPEN_CHATBOT Learners open the chatbot window.
5.4 CLOSE_CHATBOT Learners close the chatbot window.
There are two types of web pages in FLoRA, namely instructions and inspirational_materials. The former refers to the only web page detailing the assignment instruction; the latter refers to various reading materials provided by the teachers for students to get inspiration about different topics.
@@ 6.2 TABLE_OF_CONTENT Learners check the table of content, such as scrolling in that area
The writing log data contains all the keystroke events and the writing content that happen while students writing the proposal. It contains the following columns:
@@
Second, we tested the stability of the trace data across cohorts (temporal stability and replicability). We compared a pilot dataset collected under identical experimental conditions in Semester One 2024 (2024S1) with the main dataset from Semester Two 2024 (2024S2), which we intend to release as open data. After conducting normality and homogeneity‑of‑variance tests, we found that both datasets were non‑normal but exhibited equal variances. We therefore used a Mann–Whitney U test, which yielded p = 0.171063 (> 0.05), indicating no significant difference in the distribution of trace logs per student between 2024S1 and 2024S2. Figure 3 shows comparable right‑skewed distributions across semesters, with similar medians and a comparable spread of outliers; although 2024S2 includes more students, the distributional patterns remain stable. This consistency supports the reliability of the collection process and the generalisability of our findings.
@@
Fourth, we validated the assessment data (assessment reliability). Expert teachers evaluated all submitted essays using detailed, criterion‑based rubrics. Their discipline expertise and consistent application of the criteria ensured rigorous grading and supported reliable links between learning outcomes and interactions with the GenAI system.
Finally, we verified the survey data (background variable integrity). The research team reviewed all pre‑task and post‑task responses, removing incomplete entries when appropriate. This step ensured that the background information and participant feedback used in subsequent analyses were reliable.
Together, these procedures triangulate platform validity, cross‑semester stability, dialogue quality, assessment reliability, and survey integrity, thereby strengthening the overall credibility of the dataset and the conclusions drawn from it.
@@
This dataset is best suited for research about IPS, as well as for detailed writing process analytics. Researchers can use the dialogues to infer planning, monitoring, evaluation, and adaptation processes across sessions. Because the interaction log data is pre‑labelled with SRL processes in FLoRA (using Bannert’s SRL framework [31, 29]), this data can also be used for SRL analysis to evaluate students cognition and metacognition activities during the learning.

### v2

@@ Abstract
Information Problem Solving (IPS) is a critical competency for academic and professional success in education, work, and life. The advent of Generative Artificial Intelligence (GenAI), particularly tools like ChatGPT, has introduced new possibilities for supporting students in complex IPS tasks. However, empirical insights into how students engage with GenAI during IPS and how these tools can be effectively leveraged for learning remain limited. Moreover, differences in background—shaped by cultural and socioeconomic factors—pose additional challenges to the equitable integration of GenAI in educational contexts. To address this gap, we present an open-source dataset collected from 279 students at a public Australian university. The dataset was generated through students’ use of FLoRA, a GenAI-powered educational platform that is widely adopted in the field of learning analytics. Within FLoRA, students interacted with an embedded GenAI chatbot to gather information and synthesize it into data science project proposals. The dataset captures fine-grained, multi-dimensional records of GenAI-assisted IPS processes, including: (i) student–GenAI dialogue transcripts; (ii) writing process log traces; (iii) final project proposals with human-assigned assessment scores; (iv) two surveys assessing students demographic background and their prior knowledge and experience in data science and AI; and (v) surveys capturing students’ perceptions of GenAI’s effectiveness in supporting IPS and platform use experience. This dataset provides a valuable resource for advancing our understanding of GenAI’s role in educational IPS and informing the design of adaptive, inclusive AI-powered learning tools.
@@
In recent years, GenAI systems have been rapidly adopted by learners to support a wide range of educational activities, including IPS tasks [13]. Powered by large language models, GenAI tools enable learners to interact with information in conversational and iterative ways, thereby extending traditional forms of tool use in IPS. Rather than being confined to a single stage, GenAI can be integrated across multiple phases of the IPS process [44, 4]. During the early stages of IPS, learners may use GenAI to clarify task requirements, refine research questions, or decompose complex information problems into manageable sub-tasks [22, 8]. Beyond problem formulation, GenAI increasingly serves as an interactive channel for information gathering, allowing learners to obtain explanations, examples, and summaries through iterative dialogue rather than through conventional keyword-based search alone [14]. In later phases, learners may rely on GenAI to summarise retrieved materials, compare perspectives across sources, and generate integrative explanations [16], and obtain feedback on intermediate or final task products (e.g., essay drafts) to support evaluation and revision [7, 43]. Collectively, these practices indicate that GenAI functions as a flexible cognitive tool embedded throughout learners’ IPS workflows, with implications for both the nature and sequencing of IPS activities.
@@
The rapid integration of GenAI into students’ learning practices has raised important questions about its potential impact on learning processes and outcomes. While GenAI offers new opportunities to support information access, sense-making, and feedback [42], concerns have also been expressed regarding over-reliance, reduced cognitive and metacognitive engagement, and exposure to inaccuracies and biases that may compromise information quality [33, 11, 25, 5]. These competing possibilities underscore the need for systematic empirical investigation into how learners actually make use of GenAI in their everyday IPS tasks. Although a growing body of research has begun to examine students’ use of GenAI in educational contexts, existing studies have often focused on isolated IPS phases [30, 23], relied on controlled laboratory settings [34], or drawn primarily on learners’ retrospective self-reports [28], thereby providing only partial and indirect insights into real-world IPS practices [36, 24]. To address these limitations, a critical next step is to collect and openly share large-scale datasets that capture learners’ fine-grained behavioural traces as they engage in IPS processes with GenAI in authentic educational scenarios. Such datasets are essential for advancing data-driven understanding of GenAI-assisted IPS and for informing evidence-based instructional design and educational policy.
@@
To address these challenges, this paper presents a large-scale, openly available dataset capturing learners’ IPS activities in an authentic GenAI-assisted educational context. The dataset was collected through a field study involving 519 postgraduate students at a public Australian university, who completed a two-week course assignment to develop a data science project proposal. Students were supported by FLoRA, a GenAI-powered system that integrates a GPT-4o chatbot with note-taking and writing tools and is designed to provide scaffolded support rather than direct solutions [20, 19]. The dataset comprises multiple complementary data modalities, including student–GenAI dialogue logs, fine-grained interaction traces of the writing process (e.g., mouse movements, keystrokes, and clicks), the project proposals produced as task outcomes together with instructor-assigned evaluation scores, and survey data capturing learners’ background characteristics and perceptions of GenAI use. By capturing learners’ interactions with GenAI as they gather information and construct project proposals in an authentic course setting, this dataset provides a unique empirical foundation for investigating IPS processes across phases and for developing data-driven instructional and AI-based scaffolding approaches to support effective student learning.
@@ Methods
This research project rigorously adheres to the ethical standards established by Monash University. The study was approved by the Monash University Human Research Ethics Committee under Project ID 42510. The dataset was collected from students in developing a data science project proposal as an assignment in an introductory level data science course for postgraduate students. Before commencing data collection, we distributed explanatory statements and forum announcements to all students to introduce the purpose of the study and to clarify that the data collected would be used exclusively for research purposes. Given that the study required the use of FLoRA [19], we informed students that participation was optional and that they could complete the assignment without utilizing FLoRA if they preferred. We emphasised that all students had the right to withdraw from the research study at any time. Importantly, we assured them that their decision to participate or not would have no influence on their final assignment scores. By providing this information, we ensured that students could make informed decisions about their involvement, thereby upholding ethical research practices and respecting their autonomy.
@@
The participants in this study comprised 519 postgraduate students and 65 students choose not to use FLoRA. These students had enrolled in the Foundations of Data Science course, either as an elective or a core component of their respective degrees. As this research was conducted as a field study, all participants worked within the same learning environment, ensuring consistency in the educational context. Data were collected in two offerings of the same course: a pilot cohort in Semester 1, 2024 (2024S1) and the main cohort in Semester 2, 2024 (2024S2). Unless otherwise stated, the open dataset released with this paper corresponds to 2024S2; 2024S1 was used only for stability checks reported in the Technical Validation section.
@@
Data collection was conducted using the FLoRA platform [19], which offers a suite of tools tailored for educational research, including an annotation tool, an GenAI-powered chatbot tool, and a writing tool. These tools were selected to capture fine-grained information problem solving traces across key phases—students’ in-text highlighting/notes while engaging with abundant readings (annotation tool), their iterative querying and evaluation actions during AI-assisted search and reasoning (GenAI-powered chatbot), and their step-by-step proposal drafting and revision logs (writing tool). In particular, the built-in writing tool captured keystroke-level writing logs and intermediate draft states, enabling fine-grained analyses of drafting and revision as part of the synthesis/production phase of IPS. The study was systematically divided into four key phases: Pre-task survey, Training task, Main task, and Post-task survey.
Each phase was designed to collect specific types of data. FLoRA recorded fine-grained, time-stamped process data while students engaged in reading, annotating, chatting with the GenAI assistant, and writing. These data included (i) students’ annotations and notes, (ii) student–chatbot interaction logs, and (iii) time-stamped trace events generated during platform use, such as navigation events and clicks, mouse interactions (e.g., move/click/scroll), and keyboard interactions (keystrokes). Each raw event was stored as a log entry with timestamps and interface context (e.g., the active page/tool). Together, these time-stamped logs yield high-resolution, process-oriented data on learners’ information problem-solving activities.
@@ 3. Phase 3:
    GenAI-Assisted Project Proposal Task. In Phase 3, students were required to develop a data science project proposal using the FLoRA platform, supported by an embedded GenAI-powered chatbot (GPT-4o). They were first asked to select one topic from a set of 20 options spanning domains such as agriculture, education, finance, the game industry, healthcare, social media, and sports. Based on the chosen topic, students then were asked to formulate an idea for how data science could be applied in that domain and elaborate it into a written project proposal. Throughout this process, the chatbot functioned as an interactive research assistant, enabling students to ask questions, seek clarification, and obtain relevant information to support their information problem-solving activities. For example, a student exploring a topic on urban mobility could ask, “What are common data sources used in transportation studies?” or “Can you explain the concept of traffic flow prediction using machine learning?”. Also, students could use the chatbot to refine their ideas, test the relevance of their project scope, or request feedback on parts of their written proposal, such as asking, “Does this hypothesis seem focused enough?” or “Can you suggest improvements to this paragraph?” While the chatbot offered rich, responsive support, it was deliberately constrained from generating complete proposal content to ensure that students engaged meaningfully with the task and developed their own academic output. In this phase, trace data including the annotations students made, the interactions with the platform and the writing process logs was also recorded.
@@ 4. Phase 4:
In addition to these process logs, we collected students’ final proposals, rubric-based assessment scores, and survey responses as outcome and background measures.
@@
  • Missing writing trace data: Students who used external tools (e.g., Microsoft Word or Google Docs) to compose their proposals instead of FLoRA’s built-in writing tool were removed, since writing logs including keystrokes, which are crucial for tracing IPS-related behaviours, could not be captured.
@@
  • Low similarity between FLoRA and Moodle submissions: Students were required to copy their FLoRA-composed proposal into a Microsoft Word document and submit it via Moodle for grading. To ensure consistency between the analysed proposal in FLoRA and the final assessed work submitted on Moodle, we computed similarity scores between both versions. Proposals with a similarity score below 0.85 (this threshold of 0.85 was determined through manual review of the lowest 10% most dissimilar cases) were excluded due to substantial post-transfer modifications. For similarity score calculation, we utilised the cosine similarity method. Specifically, we first converted the textual data into high-dimensional dense vectors using OpenAI’s text-embedding-3-small model. This model maps text into a semantic vector space, capturing contextual meaning more effectively than traditional keyword-based methods (e.g., TF-IDF). The cosine similarity score was then computed between the resulting embedding vectors to determine the degree of semantic alignment.
@@
Students’ final project proposals were graded as part of the normal course assessment by the course teaching staff, all of whom had multiple years of teaching experience. Grading followed a criterion-based rubric. The rubric criteria included: Clear description of the goals of the project (3%), Appropriateness of topic (3%), Clear description of the business benefits and challenges (3%), Novelty/Creativity (3%), and Overall clarity of the report (3%). As this was a field study embedded in authentic teaching practice, grading was conducted following standard course procedures (i.e., no additional double-marking or sampling-based moderation was introduced solely for research purposes).
Survey responses were linked to the main dataset using students’ anonymised IDs and filtered to retain only the final analytical sample (i.e., the same 279 students retained after the interaction-log and proposal-based exclusion criteria). Responses were screened and incomplete entries for key variables were removed when appropriate. The complete survey instruments (all questions and response options) are provided in Supplementary Information Table S1, S2, and S3.
Learning Action Logs and Action Coding
FLoRA automatically parsed raw trace events into interpretable learning-action labels (Table 1) using a rule-based mapping, enabling interpretable analyses of students’ IPS behaviours [12]. Such trace data provide an unobtrusive record of learners’ moment-to-moment activities and are valuable for studying information problem solving (IPS), because IPS involves not only asking questions to the chatbot but also iterative reading, revisiting sources, writing, and tool-mediated exploration that unfold as observable interaction sequences. These traces are important for IPS research because they capture not only information seeking and evaluation behaviours, but also the synthesis/production activities (i.e., writing and revision) through which learners integrate information into an academic proposal.
In the current dataset, raw trace events were automatically mapped into learning action categories using a rule-based trace parser (i.e., an action library approach). Specifically, raw events were mapped into interpretable learning action categories (Table 1) based on the interface context and a small set of deterministic rules: (i) page context rules (e.g., events on the task-instruction interface were labelled as TASK_REQUIREMENT); (ii) reading vs. re-reading rules based on repeated visits to the same reading page (first visit labelled as READING_MATERIALS, subsequent visits labelled as REREADING_MATERIALS); (iii) navigation filtering based on dwell time (rapid page switches with dwell time << 6 seconds were labelled as PAGE_NAVIGATION); and (iv) tool- and input-based rules for writing/annotation/chatbot actions (e.g., opening/closing tools, saving, keystroke-based writing, and paste events such as PASTE_TEXT_ESSAY). Periods of inactivity longer than 5 minutes were labelled as OFF_TASK.
Importantly, we did not directly assign self-regulated learning (SRL) process labels (e.g., metacognitive planning/monitoring or cognitive reading/elaboration) in this study. Readers interested in SRL process-level coding can derive such labels by applying a process library that maps recurring sequences of learning actions to SRL processes aligned with Bannert’s model [1]. We refer readers to Fan et al. [12] for the validated action/process library framework and examples of how action sequences are mapped to SRL processes.
Table 1: The learning action categories
No. Action No. Sub_Action Sub Action Definition
0 START 0.1 START_TASK Indicate when learners start the task.
2 READING 2.1 READING_MATERIALS Learners read relevant reading materials (first visit within a session).
2.2 REREADING_MATERIALS Learners re-read relevant reading materials (revisit of a previously visited page).
3.4 SAVE_ESSAY Learners click the save button to save the essay.
3.5 CLOSE_ESSAY Learners close the essay window.
4.8 CLOSE_ANNOTATION Learners close the annotation window.
5 CHATBOT 5.1 OPEN_CHATBOT Learners open the chatbot window.
5.4 CLOSE_CHATBOT Learners close the chatbot window.
@@ 6.2 TABLE_OF_CONTENT Learners check the table of content, such as scrolling in that area
We began with the taxonomy proposed in the paper of Ren et al. [29] but found that an excessive number of student utterances were being assigned to the general-purpose category Reveal (i.e., expressing a new intent for information or refining a previous one). To enhance analytical granularity and improve interpretability, we renamed Reveal as Idea Exploration and subdivided it into six distinct student intent types: Knowledge, Judgement, Preference, Provide Context, Revise, and Inspire. The first three subcategories—Knowledge, Judgement, and Preference—were informed by Paul and Elder’s critical thinking framework [27], which distinguishes between factual, evaluative, and subjective forms of reasoning. Specifically, Knowledge utterances involve requests for definitive, factual information that is consistent across contexts (e.g., definitions or technical terms). Judgement utterances seek responses that involve critical evaluation or consideration of multiple valid perspectives. Preference utterances solicit subjective opinions or recommendations, such as which idea might be easier or more appropriate. The remaining three subcategories—Provide Context, Revise, and Inspire—emerged inductively through close examination of the collected data. Provide Context captures utterances in which students offer background information or setup without making a direct information request, often to improve the relevance of future responses. Revise applies to instances where students reformulate or elaborate on a previous question, typically in response to a chatbot answer that was unclear or insufficient. Inspire is used when students pose follow-up questions that are clearly triggered by the chatbot’s preceding response, indicating a sustained, dynamic information-seeking process.
@@
All dialogue data were annotated by three independent coders with relevant expertise in data science and educational technology. To ensure clarity and consistency in the annotation process, a detailed coding manual was developed, including operational definitions, decision rules, and illustrative examples for each code. Specifically, coders first labelled student intents and then labelled the corresponding chatbot responses. For student intents, precedence rules were applied: Inspire was used when the student explicitly extended the chatbot’s immediately preceding response (e.g., requesting more similar examples) without expressing dissatisfaction; Revise was used when the student indicated dissatisfaction with the prior response and reformulated the question, re-asked it with modifications, or added new constraints. When neither applied, coders distinguished Knowledge (requests for fixed, factual explanations/definitions) from Preference (requests for subjective opinions/recommendations, often framed as “Do you…?”) and Judgement (idea-seeking or evaluative requests with multiple valid answers, including imperatives such as “generate/give me…”). Writing-related requests were labelled as Task Clarification when asking about assignment requirements (e.g., what each section should include) and as Writing Advice when asking for editing, feedback, rewriting, or scoring of the student’s own text. Inputs that only provided isolated information or lacked a clear information need were labelled Provide Context, whereas greetings were labelled Chitchat and off-task prompts were labelled Others. For chatbot responses, coders labelled Fact for objective, well-defined answers (e.g., definitions), Suggestion for neutral guidance/options or generated content without evaluative judgement, and Critique when the response contained explicit evaluation or opinion; responses requesting additional user information were labelled Clarify, explicit inability to answer was labelled No Answer, and social/off-task replies were labelled Chitchat or Others accordingly. These rules complemented the codebook definitions and examples, ensuring consistent and reproducible labelling decisions across coders.
@@
In total, the annotated dataset comprises 3,560 interaction pairs (consisting of student prompts and corresponding chatbot responses), each labelled according to the developed codebooks described in Table 2 and Table 3, with the distribution of intent and response types visualised in Figure 1 and the distribution student activity visualised in Figure 2.
@@
The dataset is publicly available on GitHub [^1] https://github.com/Xinyu-Li/FLoRA_Open_Data_IPS under the BSD-2-Clause license. It comprises multiple CSV-format data files, as detailed below.
@@
The writing logs enable writing-process analytics (e.g., drafting/revision dynamics, writing bursts, and temporal alignment between writing and reading/chatbot use), which is central to studying IPS synthesis/production with GenAI support. The writing log data contains all the keystroke events and the writing content that happen while students writing the proposal. It contains the following columns:
@@
Second, we tested the stability of the trace data across cohorts (temporal stability and replicability). As described in the Methods, the study was conducted in two offerings of the same course: a pilot cohort in 2024S1 and the main cohort in 2024S2 (released as open data). We assessed stability by comparing the distribution of the number of time-stamped raw trace events (log entries) per student between 2024S1 and 2024S2. After conducting normality and homogeneity‑of‑variance tests, we found that both datasets were non‑normal but exhibited equal variances. We therefore used a Mann–Whitney U test, which yielded p = 0.171063 (> 0.05), indicating no significant difference in the distribution of trace logs per student between 2024S1 and 2024S2. Figure 3 shows comparable right‑skewed distributions across semesters, with similar medians and a comparable spread of outliers; although 2024S2 includes more students, the distributional patterns remain stable. This consistency supports the reliability of the collection process and the generalisability of our findings.
@@
Fourth, we verified the assessment outcome measure (rubric transparency and integrity). As described in the Methods, students’ proposals were graded by experienced course teaching staff using a criterion-based rubric with predefined criteria and weights. Using an explicit rubric provides a structured and transparent learning-outcome measure for linking performance with students’ interactions and trace data in FLoRA.
Finally, we verified survey data integrity. As described in the Methods, the research team reviewed pre-task and post-task survey responses and removed incomplete entries for key variables when appropriate. The complete survey instruments (all questions and response options) are provided in Supplementary Information Table S1, S2 and S3, enabling replication and reuse.
Together, these procedures triangulate platform validity, cross-semester stability of trace data, dialogue content integrity, rubric-based outcome transparency, and survey data integrity, thereby strengthening the overall credibility of the dataset and the conclusions drawn from it.
@@
This dataset is best suited for research about IPS, including fine-grained analysis of students’ writing and revision processes during the synthesis/production stage while interacting with GenAI. Researchers can use the dialogue sequences—together with behavioural and writing logs—to examine indicators of planning, monitoring, evaluation, and adaptation across sessions. Because the interaction log data is pre-labelled with learning-action categories (Table 1), it can be used to analyse students’ IPS behaviours at a fine temporal granularity. Researchers interested in SRL process-level analysis can further map sequences of learning actions to Bannert-aligned cognitive and metacognitive SRL processes by applying an established process library (see Fan et al. [12]).